Nine cards are marked with the numbers 1,2,3,4,5,6,7,8, and 9 then shuffled, and the three cards are drawn. How many different three-card combinations are possible? How many three-card hands contain a number less than three?
Answer by stanbon(75887)

Nine cards are marked with the numbers 1,2,3,4,5,6,7,8, and 9 then shuffled, and the three cards are drawn.
How many different three-card combinations are possible?
Ans: 9C3 = (9*8*7)/(1*2*3) = 84
---------------------------------------
How many three-card hands contain a number less than three?
# containing only numbers greater than 3 is 5C3 = 10
----
Therefore # containing at least one number less than 3 is 84-10 = 74
=============================================
Cheers,
Stan H.